find an equation of a line in standard form that has x intercept 6 and y intercept 4
Found 2 solutions by richwmiller, mananth:
Answer by richwmiller(17219)

x intercept means y=0 and here x=6 point (6,0)
y intercept means x=0 and here y=4 point (0,4)
find the equation of a line with these two points.
I hope you see that mananth's answer is wrong.
y=mx+b
m=-2/3 is correct but we already know that y=4 is the y intercept not -4
so
y=-2x/3+4 not -4

Answer by mananth(16949)


x intercept =6 point = (6,0)
y intercept = 4 point = (0,4)
x1 y1 x2 y2
6 0 0 4

slope m =(y2-y1)/(x2-x1)
(4-0)/( 0-6)
(4/-6)
m=- 2/ 3

Plug value of the slope and point(6,0 ) in
Y = m x + b
0=-4 + b
b= 0 - 4
b= -4
So the equation will be
Y =(- 2/3)x -4
m.ananth@hotmail.ca